# AI Should Sense Better, Not Just Scale Bigger: Adaptive Sensing as a Paradigm Shift

### v2

@@ 1 Introduction
However, current AI methodologies still primarily address distribution shift by scaling neural architectures and datasets oquab2023dinov2 ; cherti2023openclip ; achiam2023gpt ; radford2023robust , leaving the sensing interface unchanged. A few existing sensor-aware strategies, such as active perception, which repositions sensors or robots to optimize viewpoints bajcsy1988active ; shi2023real , and sensor-fusion budgeting, which schedules when and what modalities to activate tilmon2023energy , operate at the system or motion-planning level. They neglect how raw analog signals are digitized in the first place to best serve the downstream model. Only recently have benchmarks begun isolating the impact of intrinsic sensor settings, such as exposure and gain, on recognition accuracy baek2024unexplored ; baek2025adaptive . Inspired by these insights, the concept of test-time input adaptation has emerged: dynamically optimizing sensor parameters frame-by-frame to deliver model-friendly data. A first prototype, Lens baek2025adaptive , significantly outperforms traditional (human-oriented) auto-exposure methods, maintaining high classification accuracy despite a 50$\times$ reduction in model size. Preliminary evidence suggests even greater potential: with ideal sensor adaptation, a lightweight 5M-parameter EfficientNet-B0 manion2024effect can surpass the 632M-parameter OpenCLIP-H openclip trained on 160$\times$ more training data baek2024unexplored . These early results highlight both the promise of adaptive sensing and the urgency for systematic exploration.
@@
Recent embodied-AI benchmarks in locomotion sferrazza2024humanoidbench ; li2023behavior , household manipulation srivastava2022behavior , and interactive task execution yang2025embodiedbench exemplify an urgent reality, where the future AI must sense, reason, and interact from and with the real world within strict real-time and on-device constraints. Physical AI platforms from micro-drones to neural prostheses face inherent limits in size, energy, and thermal dissipation, suggesting limits to the traditional “bigger model, bigger dataset” paradigm. Adaptive sensing provides a complementary solution: by reshaping inputs directly at the source, it delivers model-friendly signals that enable smaller networks to thrive. Just as the human visual system couples rapid eye movements to visual cognition, next-generation AI systems must integrate closed-loop, real-time sensor control and preprocessing directly into their inference pipeline.
@@ 2 Current State of AI: Limitations of the Model-Centric Paradigm
  • Economic Sustainability and Scalability: The continuous increase in model complexity and training data volumes imposes a substantial financial burden strubell2020energy ; castro2024artificial . The financial costs associated with advanced computational infrastructure, specialized hardware, and extensive data acquisition quickly become prohibitive. This economic barrier hampers the scalability and widespread adoption of advanced AI technologies, especially in resource-constrained settings or smaller-scale enterprises.
@@ 3 Adaptive Sensing as a Necessary Paradigm Shift
Vision Pupil diameter 2–8 mm ($\sim$16$\times$ light gain) in < 200 ms. Fixed or two-step aperture; ISO/shutter in coarse steps gutschick2002should .
Dark adaptation restores sensitivity; ciliary muscles refocus 10 cm to infinity. Autofocus 50–300 ms; quantum efficiency and CFA static.
@@
3.1 Early Evidence from Image Classification
Recent work on ImageNet-ES baek2024unexplored and ImageNet-ES-Diverse baek2025adaptive evaluates real-world covariate shifts induced by controlled sensor variation. Lens baek2025adaptive (Figure 1) is the first model-friendly, test-time input adaptation framework to mitigate covariate shifts. It operates in a post-hoc, adaptive, and camera-agnostic sensor control manner, dynamically responding to scene characteristics based on VisiT scores to provide optimal image quality for neural networks. The key empirical findings are as follows, which underscore both the promise and the necessity of adaptive sensing methodologies:
  • Accuracy gain: Adaptive sensing significantly improves accuracy up to 47.58%p without model modification or maintain accuracy despite 50$\times$ model size difference.
@@ 3.2 Advantages for Real-World Agents and Deployment
  • Broad Domain Applicability: Adaptive sensing has the potential to practically impact diverse domains that utilize sensor data. Concrete scenarios are as follows:
    • Humanoid Robotics: Adaptive multimodal sensor adjustments—such as dynamically modulating visual, auditory, and tactile sensors—to improve real-time balance control, manipulation tasks, and interaction quality in complex, unstructured environments.
    • Healthcare Diagnostics: Dynamic adjustment of medical imaging parameters to optimize image quality based on patient-specific characteristics and clinical contexts (e.g., adjusting MRI sequences dynamically to enhance diagnostic accuracy).
    • Autonomous Vehicles: Real-time optimization of camera and lidar parameters to rapidly adapt to changing lighting conditions and weather (e.g., automatic adjustments during fog, rain, or sudden brightness changes to enhance road safety).
    • Agriculture Monitoring: Adaptive drone sensing systems dynamically adjusting sensor settings to capture high-quality data under varying conditions, such as different crop types, growth stages, or environmental stress factors (e.g., real-time optimization for precise crop health monitoring).
    • Environmental Monitoring: Adaptive sensing to optimize sensor settings dynamically for capturing accurate data on air and water quality under varying environmental conditions, thereby improving monitoring accuracy and predictive modeling.
@@ 5 Closed-Loop Framework for Embodied AI Agents: Towards Humanoid
(b) Single-Shot-Perception (Lens baek2025adaptive ).
@@
To fill the gaps and integrate adaptive sensing into embodied reinforcement learning (RL), we propose a principled Adaptive Sensing Framework, starting from the most fundamental RL setting, a standard MDP formulation puterman2014markov , and progressively extending it to accommodate single-shot perception, continuous perception tasks, sensorimotor interactions, and ultimately multimodal embodied scenarios, all within a fully closed-loop adaptive structure. The ultimate goal is to establish a robust, adaptive AI pipeline capable of efficiently learning and reliably adapting to dynamic environments.
@@ 5.1.1 Common Notations
Single-Shot Adaptive Sensing (Perception-Only, No Actions, (Fig. 3(b))). We define a newly augmented stochastic process $\mathcal{P}:=(\mathcal{S},\mathcal{O},P^{\prime}_{E},Q_{M})$ defined by $P^{\prime}_{E}$ which is the environment transition probability similar to $P_{E}$ defined in Eq.(1), but agnostic to actions. This new stochastic process $\mathcal{P}$ captures ways in which sensing configurations are adaptively selected per-timestep for single-shot perception tasks (e.g., image classification tasks).
Concretely, at timestep $t$, the agent performs:
@@ 6 Challenges and Counterarguments
While adaptive sensing offers transformative potential to reshape AI design and operation, its widespread adoption faces several critical challenges:
@@
  • Performance Trade-offs: Adaptive sensing strategies may initially not match peak performance of large-scale, resource-intensive models. Stakeholders may prioritize immediate maximal accuracy over long-term sustainability and efficiency.
@@ 7 Open Research Directions
  • Algorithmic Innovation in Real-time Adaptation: Advance efficient algorithms using reinforcement learning sutton1998reinforcement , adaptive control landau2011adaptive , and online learning hoi2021online to optimize sensor parameters in dynamic environments. Developing frameworks to explore and exploit sensor configuration spaces will enhance responsiveness and efficiency.
@@ 8 Conclusion
We have identified critical gaps in existing methodologies, their limited consideration of continuous, closed-loop interactions essential for real-world embodied agents. Addressing these gaps through targeted research—such as standardized adaptive benchmarks, real-time adaptation algorithms, privacy and ethical standards, and multimodal sensor integration—will be vital. There are urgent, interdisciplinary opportunities that the broader AI community is uniquely positioned to tackle.
Given the escalating environmental, economic, and societal pressures resulting from conventional AI scaling, embracing adaptive sensing is not simply advantageous; it is imperative. We call upon the AI community to prioritize this research agenda. Now is the moment for decisive collective action—by integrating adaptive sensing, we can ensure AI’s trajectory aligns fundamentally with ecological responsibility, ethical integrity, and global equity.

### v3

@@ 1 Introduction
However, current AI methodologies still primarily address distribution shift by scaling neural architectures and datasets oquab2023dinov2; cherti2023openclip; achiam2023gpt; radford2023robust, leaving the sensing interface unchanged. A few existing sensor-aware strategies, such as active perception, which repositions sensors or robots to optimize viewpoints bajcsy1988active; shi2023real, and sensor-fusion budgeting, which schedules when and what modalities to activate tilmon2023energy, operate at the system or motion-planning level. Post-hoc physics-based agarwal2021simulation; planche2021physics and physics-informed (e.g., PINNs, DSE) raissi2019pinns; maier2018dse sensor simulations cannot recover the coupled environment–sensor dynamics once analog signals are digitized, as the measurements have already been shaped by sensor configurations, leading to irreversible information loss. They neglect how raw analog signals are digitized in the first place to best serve the downstream model. Historically, adaptive optics (e.g., radar CFAR finn1966radar; rohling2007radar, astronomical seeing correction babcock1953astronomical) were designed to enhance human perception or measurement fidelity, but were grounded in physical or human-centric criteria rather than how sensors adapt to model’s representation space.
[…]
@@
Recent embodied-AI benchmarks in locomotion sferrazza2024humanoidbench; li2023behavior, household manipulation srivastava2022behavior, and interactive task execution yang2025embodiedbench exemplify an urgent reality, where the future AI must sense, reason, and interact from and with the real world within strict real-time and on-device constraints. Physical AI platforms from micro-drones to neural prostheses face inherent limits in size, energy, and thermal dissipation, suggesting limits to the traditional “bigger model, bigger dataset” paradigm. Adaptive sensing provides a complementary solution: by reshaping inputs directly at the source (i.e., at the hardware level), it delivers model-friendly signals that enable smaller networks to thrive. Just as the human visual system couples rapid eye movements to visual cognition, next-generation AI systems must integrate closed-loop, real-time sensor control and preprocessing into their inference pipeline.
@@ 2 Current State of AI: Limitations of the Model-Centric Paradigm
  • Economic Sustainability and Scalability: The continuous increase in model complexity and training data volumes imposes a substantial financial burden strubell2020energy; castro2024artificial. The costs associated with advanced computational infrastructure, specialized hardware, and extensive data acquisition quickly become prohibitive. This economic barrier hampers the scalability and widespread adoption of advanced AI technologies, especially in resource-constrained settings or smaller-scale enterprises.
@@ 3 Adaptive Sensing as a Necessary Paradigm Shift
[…]
Dark adaptation restores sensitivity; ciliary muscles refocus from 10 cm to infinity. Autofocus 50–300 ms; quantum efficiency and CFA static.
@@
3.1 Early Evidence from Vision Tasks
Recent work on ImageNet-ES baek2024unexplored and ImageNet-ES-Diverse baek2025adaptive evaluates real-world covariate shifts induced by controlled sensor variation. Lens baek2025adaptive (Figure 1) is the first model-friendly, test-time input adaptation framework to mitigate covariate shifts for image classification. It operates in a post-hoc, adaptive, and camera-agnostic sensor control manner, dynamically responding to scene characteristics based on VisiT scores to provide optimal image quality for neural networks. The key empirical findings are as follows, which underscore both the promise and the necessity of adaptive sensing methodologies:
[…]
@@ 3.1 Early Evidence from Vision Tasks
Beyond image classification, adaptive sensing has shown consistent benefits across diverse vision tasks. In more complex settings such as object detection, segmentation, and remote photoplethysmography (rPPG), adaptive control of sensor parameters or viewpoints improves model robustness and accuracy under environmental shifts compared to non-adaptive configurations (e.g., fixed or auto-exposure, static viewpoints) baek2025adaptive; madeye; odinaev2023rPPG. In 6D object-pose estimation—a multimodal task involving both depth and RGB channels—adaptive cross-modal coordination marks a new stage of adaptive sensing by jointly optimizing multiple sensing modalities han2025senseshift6d. It shows that the multi-modal control can achieve higher accuracy and stability than single-modal controls (e.g., RGB-only or depth-only) and surpasses factory defaults, while further enhancing robustness and data efficiency. Together, these findings establish adaptive sensing as a scalable, complementary, and generalizable paradigm for robust perception across diverse sensing modalities and task conditions.
@@ 3.2 Advantages for Real-World Agents and Deployment
Table 2: Broad real-world applicability of adaptive sensing.
Domain Role of Adaptive Sensing
Humanoids Adaptive multimodal sensor adjustments—such as dynamically modulating visual, auditory, and tactile sensors—improve real-time balance, manipulation, and interaction quality in complex, unstructured environments.
Healthcare Dynamic adjustment of medical imaging parameters optimizes image quality based on patient-specific characteristics and contexts (e.g., dynamically adjusting MRI sequences to enhance diagnostic accuracy).
Self-driving Real-time optimization of camera and lidar parameters allows rapid adaptation to changing lighting and weather (e.g., fog, rain, or sudden brightness changes), enhancing perception robustness and road safety.
Agriculture Adaptive drone sensing systems dynamically adjust settings to capture high-quality data under varying conditions—such as crop type, growth stage, and environmental stress—for precise crop health monitoring.
Environment Adaptive sensing dynamically tunes sensor settings to capture accurate air and water quality data under diverse environmental conditions, improving monitoring accuracy and predictive modeling.
@@
  • Broad Domain Applicability: Adaptive sensing has the potential to practically impact diverse domains that utilize sensor data. Concrete scenarios are provided in Table 2.
@@ 5 Closed-Loop Framework for Embodied AI Agents: Towards Humanoid
(b) Single-Shot-Perception (non-RL, Lens baek2025adaptive).
@@
To fill the gaps and integrate adaptive sensing into embodied AI across both reinforcement learning (RL) and non-RL settings, we propose a principled Adaptive Sensing Framework. The framework is scalable and covers essential settings for embodied AI, starting from the most fundamental RL setting, a standard MDP formulation puterman2014markov as sequential decision-making. It also encompasses non-RL settings, drawing on single-shot control scenarios explored in early adaptive sensing studies, and progressively extends tocontinuous perception tasks, sensorimotor interactions, and ultimately multimodal embodied scenarios, all within a fully closed-loop adaptive structure. The ultimate goal is to establish a robust, adaptive AI pipeline capable of efficiently learning and reliably adapting to dynamic environments.
@@ 5.1.1 Common Notations
Single‐Shot Adaptive Sensing (Perception-Only, No Actions, (Fig. 3(b))). We define a newly augmented stochastic process $\mathcal{P}:=(\mathcal{S},\mathcal{O},P^{\prime}_{E},Q_{M})$ defined by $P^{\prime}_{E}$ which is the environment transition probability similar to $P_{E}$ defined in Eq.(1), but agnostic to actions. This new stochastic process $\mathcal{P}$ captures ways in which sensing configurations are adaptively selected per-timestep for single-shot perception tasks (e.g., image classification tasks). Concretely, at timestep $t$, the agent performs:
@@ 6 Challenges and Counterarguments
While adaptive sensing offers transformative potential to reshape AI design and operation, its widespread adoption faces several critical challenges and counterarguments.
6.1 Challenges
@@
  • Performance Trade-offs: Adaptive sensing strategies may at first underperform large-scale, resource-intensive models and may face real-time or deployment bottlenecks in practice. Stakeholders may prioritize immediate maximal accuracy over long-term sustainability and efficiency.
@@
6.2 Counterarguments
  • Proprietary Sensors & Transparency: Although many sensors are proprietary, adaptive sensing does not require internal hardware access; existing APIs for exposure or gain control are sufficient. As demand grows, vendors are likely to expand such interfaces without compromising intellectual property. Overall, adaptive sensing is more likely to increase, rather than reduce, transparency.
  • Alternative methods: Approaches such as efficient architectures and federated learning (FL), often proposed within the current model-centric paradigm to address generalization and computational challenges, cannot recover information lost at capture time or resolve environment–sensor shifts. They also introduce trade-offs—efficiency may reduce OOD robustness and fairness, while FL suffers from non-IID bias and privacy risks. Adaptive sensing complements these methods by operating upstream, allowing compact models to match or even surpass large-scale ones.
  • Task-dependent or rapidly changing configurations: In single-task deployments, Lens selects from a small runtime candidate pool, so task dependence is not limiting. For multi-tasking, condition the pool on the active task with a conditional/hierarchical policy. Under dynamics, avoid per-frame updates by triggering switches on drift/performance proxies (e.g., confidence, calibration, quality drops), stabilize with hysteresis and switch-cost/latency budgets, and schedule changes via history-/resource-aware bandits or reinforcement learning.
  • When and Why Adaptive Sensing Excels: Adaptive sensing outperforms domain adaptation, domain generalization, and test-time adaptation under covariate shifts, as hardware-level control directly reshapes raw measurements and can preserves information that post-hoc model adaptation cannot recover (§1). In contrast, under semantic or content shifts (e.g., unseen classes), model-level adaptation remains more effective. The two paradigms are complementary, addressing orthogonal failure modes and achieving the best performance when combined. In resource-constrained settings, adaptive sensing remains attractive for its efficiency, enabling small models to rival much larger ones (§3). Determining when to apply each approach remains an open question (§7), as current OOD and uncertainty detectors primarily flag semantic shifts, leaving reliable differentiation between covariate and semantic (or mixed) shifts underexplored.
  • Why Closed-Loop? (vs. Separated): A standalone controller (as in Lens baek2025adaptive) is feasible and effective in low-dynamic environments where no action policy is involved (Fig. 3(b)). However, in embodied AI settings characterized by dynamic covariate shifts and sparse rewards, an optimal sensing strategy must be carefully designed to provide richer feedback and mitigate reward sparsity, as it depends on both the model’s perception characteristics and task objectives. Closed-loop co-optimization aligns sensor control with the agent’s policy learning, stabilizing observations and improving both sample efficiency and robustness (§4–5).
@@ 7 Open Research Directions
  • Algorithmic Innovation in Real-time Adaptation: Advance efficient algorithms using reinforcement learning sutton1998reinforcement, adaptive control landau2011adaptive, and online learning hoi2021online to optimize sensor parameters in dynamic environments. Developing frameworks to explore and exploit sensor configuration spaces will enhance responsiveness and efficiency. Future directions also include addressing practical deployment challenges such as multi-sensor synchronization, single-sensor multitasking, and deadline-constrained adaptation, which are key to mitigating real-time bottlenecks.
@@ 8 Conclusion
We have identified critical gaps in existing methodologies, their limited consideration of continuous, closed-loop interactions essential for real-world embodied agents. Addressing these gaps through targeted research—such as standardized adaptive benchmarks, real-time adaptation algorithms, privacy and ethical standards, and multimodal sensor integration—will be vital. There are urgent, interdisciplinary opportunities that the broader NeurIPS community is uniquely positioned to tackle.
Given the escalating environmental, economic, and societal pressures resulting from conventional AI scaling, embracing adaptive sensing is not simply advantageous; it is imperative. We call upon the NeurIPS community to prioritize this research agenda. Now is the moment for decisive collective action—by integrating adaptive sensing, we can ensure AI’s trajectory aligns fundamentally with ecological responsibility, ethical integrity, and global equity.
Acknowledgments
This work was supported in part by Institute of Information communications Technology Planning Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02263754), in part by the National Research Foundation (NRF) of Korea grant funded by the Korea government (MSIT) (No. RS-2023-00222663).